OpenStudy (anonymous):

Joyce needs to gather data that can be modeled with a linear function. Which situation would give Joyce the data she needs? A. the area of a square and the length of its side B. the area of a circle and the length of the its radius C. the perimeter of a square and the length of its side D. the volume of a cylinder and the length of its radius

OpenStudy (anonymous):

joyce should get back in the kitchen

OpenStudy (anonymous):

that steak isn't gonna cook itself

OpenStudy (tkhunny):

Square: \(Area = Side^{2}\) That's not linear. Circle: \(Area = \pi\cdot Radius^{2}\) That's not linear. Square: \(Perimeter = 4\cdot Side\) That IS linear. Cyllider: \(Area = \pi\cdot Radius^{2}\cdot Height\) That's not linear

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

!1111!!1!11!!!!!!!1111